# VoiceSculptor: Your Voice, Designed By You

### v1

@@ Abstract
Despite rapid progress in text-to-speech (TTS), open-source systems still lack truly instruction-following, fine-grained control over core speech attributes (e.g., pitch, speaking rate, age, emotion, and style). We present VoiceSculptor, an open-source unified system that bridges this gap by integrating instruction-based voice design and high-fidelity voice cloning in a single framework. It generates controllable speaker timbre directly from natural-language descriptions, supports iterative refinement via Retrieval-Augmented Generation (RAG), and provides attribute-level edits across multiple dimensions. The designed voice is then rendered into a prompt waveform and fed into a cloning model to enable high-fidelity timbre transfer for downstream speech synthesis. VoiceSculptor achieves open-source state-of-the-art (SOTA) on InstructTTSEval-Zh, and is fully open-sourced, including code and pretrained models, to advance reproducible instruction-controlled TTS research.
@@ 1 Introduction
In recent years, the rapid evolution of large-scale multimodal foundation models has fundamentally reshaped the paradigm of generative artificial intelligence (AI), enabling unified generation across text, speech, images, and videos. Commercial systems such as Gemini 2.5 Pro & Flash, and GPT-4o mini have demonstrated strong instruction-following and multimodal reasoning capabilities, while recent text-to-video and audio–visual generation models, including Veo 3, Wan 2.6, Seedance 1.5 pro and Kling 2.6, have shown impressive progress in jointly synthesizing coherent visual content and synchronized audio. These advances highlight a growing trend toward holistic, natural-language-driven content creation, in which users expect to express complex intentions once and obtain rich and accurate multi-modal outputs.
@@
In parallel, audio-centric foundation models such as MiMo-Audio (Zhang et al., 2025a) and Step-Audio2 (Wu et al., 2025) have significantly improved the representation and generalization capacity of speech models, laying the foundation for unified modeling of text and audio, and enabling speech generation that can be directly controlled by textual instructions. Modern neural text-to-speech (TTS) systems, including CosyVoice2 (Du et al., 2024) , LLaSA (Ye et al., 2025) , F5-TTS (Chen et al., 2025) , SparkTTS (Wang et al., 2025) , and Index-TTS2 (Zhou et al., 2025) , can now generate highly natural speech and effectively mimic speaker timbre when reference audio is available. However, despite these advances, controllability over speech attributes remains limited, especially when compared to the flexibility observed in recent multimodal generation systems.
Traditional non-LLM-based instruct TTS approaches, such as PromptStyle (Liu et al., 2023) and PromptSpeaker (Zhang et al., 2023) , rely on predefined prompts or learned style or speaker embeddings to modulate speech attributes. While effective in constrained scenarios, these methods suffer from limited scalability and expressiveness, as they depend on carefully designed templates and struggle to generalize to open-ended natural language instructions. Recent LLM-based approaches attempt to leverage the semantic modeling capability of large language models for instruction-driven speech synthesis. VoxInstruct (Zhou et al., 2024) and FleSpeech (Li et al., 2025) represent this emerging direction by conditioning speech generation on textual or multimodal representations. In particular, FleSpeech introduces a multimodal prompt encoder and aligns text, audio, and visual representations into a shared latent space to guide diffusion-based speech generation. Similarly, HiStyle (Zhang et al., 2025c) adopts a representation alignment strategy by mapping text embeddings, style embeddings, and speaker embeddings into a unified embedding space, enabling textual descriptions to implicitly steer the synthesis process. By aligning heterogeneous modalities or control signals, these methods improve the consistency between input instructions and generated speech styles. However, despite their effectiveness, both FleSpeech and HiStyle rely on compact, continuous embeddings as the primary control interface. Such embedding-level conditioning inevitably compresses rich and multi-dimensional voice attributes into low-bandwidth representations, limiting the ability of large language models to perform explicit reasoning over fine-grained acoustic properties. As a result, controllability remains coarse and implicit, making it difficult to precisely manipulate individual attributes or faithfully interpret complex, compositional natural language instructions.
@@
Nevertheless, most existing systems still generate speech primarily conditioned on text and reference audio, offering limited direct control over fine-grained acoustic attributes such as pitch, speaking rate, age, emotional expression, speaking style, and others. This gap reveals a fundamental bottleneck in current generative systems: although natural language has become the dominant interface for controlling complex multimodal generation, speech synthesis still lacks a principled and flexible mechanism to translate high-level linguistic intent into fine-grained acoustic realization. In contrast to visual and video generation, voice generation remains heavily constrained by reference-based conditioning or rigid control tokens.
@@
To address this limitation, we propose VoiceSculptor, a unified and highly flexible speech synthesis framework that bridges natural language intent and fine-grained voice generation. Unlike conventional TTS systems that rely on fixed control tokens or reference audio alone, VoiceSculptor enables users to design speaker timbre and manipulate multiple voice attributes directly via free-form natural language instructions.
@@ 2.1 Overview
As shown in Figure 1, VoiceSculptor adopts LLaSA-3B (Ye et al., 2025) as the voice design model and CosyVoice2 (Du et al., 2024) as the voice cloning model. LLaSA is built upon the open-source LLaMA (Touvron et al., 2023) family released by Meta and is fine-tuned to leverage the strong text understanding and sequence modeling capabilities of large language models. To enable speech generation within an LLM framework, LLaSA incorporates an advanced neural audio codec, XCodec2, which converts continuous speech waveforms into discrete audio tokens that resemble text tokens.
With the introduction of XCodec2, speech synthesis is reformulated as a sequence-to-sequence generation problem. Given a natural language instruction, the LLM is responsible for interpreting the semantic and stylistic intent of the text and predicting a corresponding sequence of discrete audio tokens. These predicted tokens are then decoded by the audio codec to reconstruct high-quality speech waveforms. This design allows the model to jointly reason over text and speech tokens using a unified autoregressive generation paradigm.
Built upon this foundation, VoiceSculptor consists of three key contributions. First, we construct a comprehensive data processing pipeline that supports large-scale data collection, filtering, multi-dimensional annotation, and human verification, providing high-quality supervision for instruction-driven voice design. Second, we introduce CoT–based fine-grained attribute modeling, which enables precise and interpretable control over prosodic and stylistic attributes by guiding the LLM to reason explicitly over attribute-related semantics during audio token generation. Third, to improve robustness and generalization to diverse and out-of-domain natural language instructions, we incorporate a RAG mechanism that supplies semantically aligned in-domain instruction examples at inference time, effectively grounding the instruction interpretation process.
Together, these components enable VoiceSculptor to translate natural language instructions into controllable and reusable voice representations, which can be seamlessly consumed by downstream speech synthesis models such as CosyVoice2.
@@ 2.2 Data Processing Pipeline
For linguistic alignment, automatic speech recognition (ASR) and forced alignment are applied to obtain accurate transcriptions and punctuation information. Specifically, FireRedASR (Xu et al., 2025b) is used to transcribe Chinese audio, Whisper (Radford et al., 2022) is employed for English audio, and SenseVoice (An et al., 2024) is used for ASR cross-validation, as well as language and emotion recognition.The resulting transcripts are further processed using the Kaidi  [^5] https://github.com/sunsetsonwheels/kaidi alignment tool to perform character-level (for Chinese) or word-level forced alignment, producing precise timestamps and pause durations between adjacent tokens. Based on these fine-grained temporal cues, a trained punctuation prediction model is applied to restore punctuation marks, yielding linguistically coherent and temporally aligned text–audio pairs.
@@
Based on the aligned text–audio representations, a structured annotation process is subsequently conducted across multiple levels. At the advanced acoustic level, each audio sample is first analyzed using Gemini 2.5 Pro to obtain multi-dimensional annotations, including pitch, speaking rate, loudness, speaker gender and age, emotional state, paralinguistic characteristics, and contextual attributes. Based on these structured annotations, DeepSeek is subsequently employed to perform translation and caption generation, producing natural-language descriptions of the audio style and vocal characteristics. To mitigate hallucinations introduced by large language models, we apply a set of rule-based regular expression filters to remove inconsistent or unsupported content. This process yields relatively high-quality, semantically grounded annotations and captions that serve as reliable supervision for instruction-driven voice design. At the emotional level, speech is annotated through a cross-validated emotion labeling process. Specifically, we employ multiple complementary models, including Emo2Vec (Ma et al., 2024) , Qwen3-8B (Yang et al., 2025) , SenseVoice, and Qwen3-Omni (Xu et al., 2025a), to independently predict emotion-related attributes from the audio. The outputs from these models are then cross-validated to resolve inconsistencies and improve label reliability, resulting in a final set of emotion annotations with higher robustness and accuracy. At the prosodic level, fine-grained acoustic attributes are annotated through a combination of automatic estimation, statistical analysis, and human verification. Specifically, we employ the DataSpeech (Lyth and King, 2024) model to estimate continuous prosodic features, including the mean and standard deviation of pitch, mean and standard deviation of energy, speaking rate, and loudness. Speaker gender and age are annotated using VoxProfile (Feng et al., 2025) . Based on the extracted attributes, we perform distributional analysis across the dataset and conduct targeted human listening to calibrate attribute boundaries. The continuous prosodic features are discretized into five intervals for each dimension, while age is categorized into four groups (child, youth, middle-aged, and elderly). Finally, for samples originating from in-house datasets, we leverage available ground-truth annotations to correct and refine the predicted age and gender labels, resulting in a set of reliable and structured prosodic annotations.
@@
Through this comprehensive and iterative pipeline, VoiceSculptor constructs a high-quality, multi-dimensionally annotated dataset that supports natural-language-driven voice control, fine-grained attribute manipulation, and robust instruction understanding.
@@ 2.3 CoT–based Fine-grained Attribute Modeling
Various fine-grained acoustic attributes in speech signals—such as pitch, loudness, speaking rate, and temporal dynamics—play a critical role in shaping prosodic structure, expressive rhythm, and overall vocal style. However, directly conditioning speech synthesis models on explicit attribute tokens often leads to brittle control and over-reliance on structured inputs, limiting the model’s ability to generalize to diverse natural language instructions.
@@
Furthermore, to prevent over-dependence on explicit attribute tokens and to encourage deeper instruction understanding, we introduce a stochastic attribute token dropout strategy during training. Fine-grained attribute tokens are randomly removed from the input with a predefined probability such as 0.2, forcing the model to infer the intended acoustic attributes from natural language instructions and contextual cues alone. This training strategy acts as an effective regularizer and improves the model’s robustness and generalization in instruction-driven voice control.
@@ 2.4 Retrieval-augmented Instruction Generalization
During inference, when retrieval augmentation is enabled, the incoming natural language instruction is first converted into a dense vector representation using the same embedding model. A cosine similarity–based semantic search is then performed against the vector database to identify the most relevant in-domain instructions. The retrieved instructions, which are semantically aligned with the input query, are subsequently injected into the model input, guiding the model toward more stable and accurate interpretation of the user’s intent.
@@ 3 Experiments
We conduct extensive experiments to comprehensively evaluate the effectiveness, controllability, and scalability of the proposed VoiceSculptor framework, with a particular focus on its voice design (VD) module. Given that our training data primarily consists of Chinese instruction–speech pairs, we adopt the InstructTTSEval-Zh benchmark (Huang et al., 2025) as the main evaluation protocol to assess instruction-following performance in a controlled and fair setting. We compare VoiceSculptor against both strong open-source baselines and representative commercial systems, and further analyze its performance under different model sizes, data scales, and training strategies. In addition, we perform a series of ablation studies to isolate the contributions of key components, including CoT-based fine-grained attribute tokens, text-side cross-entropy supervision, and RAG. Human subjective evaluations are also conducted to complement automated metrics and validate perceptual instruction adherence. Overall, the experimental results consistently demonstrate that VoiceSculptor achieves state-of-the-art performance among open instruction-following TTS systems, while exhibiting robust scalability and strong controllability across a wide range of settings.

### v2

@@ Abstract
Despite rapid progress in text-to-speech (TTS), open-source systems still lack truly instruction-following, fine-grained control over core speech attributes (e.g., pitch, speaking rate, age, emotion, style, and others). We present VoiceSculptor, an open-source unified system that bridges this gap by integrating instruction-based voice design and high-fidelity voice cloning in a single framework. It generates controllable speaker timbre directly from natural-language descriptions, supports iterative refinement via Retrieval-Augmented Generation (RAG), and provides attribute-level edits across multiple dimensions. The designed voice is then rendered into a prompt waveform and fed into a cloning model to enable high-fidelity timbre transfer for downstream speech synthesis. VoiceSculptor achieves open-source state-of-the-art (SOTA) on InstructTTSEval-Zh, and is fully open-sourced, including code and pretrained models, to advance reproducible instruction-controlled TTS research.
@@ 1 Introduction
In recent years, the rapid evolution of large-scale multimodal foundation models has fundamentally reshaped the paradigm of generative artificial intelligence (AI), enabling unified generation across multiple modalities, including text, speech, images, and videos. Commercial systems such as Gemini 2.5 Pro, Gemini 2.5 Flash (Comanici et al., 2025) and GPT-4o mini  [^1] https://platform.openai.com/docs/models/gpt-4o-mini have demonstrated strong instruction-following and multimodal reasoning capabilities. In contrast, recent text-to-video and audio–visual generation models, including Veo 3  [^2] https://storage.googleapis.com/deepmind-media/veo/Veo-3-Tech-Report.pdf, Wan 2.6  [^3] https://www.wan-ai.co/wan-2-6, Seedance 1.5 Pro (Seedance et al., 2025), and Kling 2.6  [^4] https://app.klingai.com/global/, have shown impressive progress in jointly synthesizing coherent visual content with synchronized audio. Text-driven generation has become the dominant interface across modalities, with text-to-image, text-to-video, and video–audio joint generation systems all supporting natural-language descriptions as the primary control mechanism.
@@
At the same time, prior speech generation research has primarily focused on zero-shot voice synthesis, while paying comparatively less attention to fine-grained, instruction-driven, controllable speech generation from natural language. Modern neural text-to-speech (TTS) systems, including CosyVoice2 (Du et al., 2024), LLaSA (Ye et al., 2025), F5-TTS (Chen et al., 2025), SparkTTS (Wang et al., 2025), and Index-TTS2 (Zhou et al., 2025), can now generate highly natural speech and effectively mimic speaker timbre when reference audio is available. However, despite these advances, controllability over speech attributes remains limited, especially when compared to the flexibility observed in recent multimodal generation systems.
Beyond zero-shot capability, recent studies are increasingly exploring fine-grained, editable, and instructi-
on-conditioned speech generation, aiming to provide users with more precise and expressive control over generated content across speech domains. To better understand this emerging line of research, it is essential to review how controllability has traditionally been achieved in speech synthesis.
Earlier controllable TTS methods typically relied on explicit prompt engineering or learned latent representations, rather than directly modeling instruction semantics from natural language, such as PromptStyle (Liu et al., 2023) and PromptSpeaker (Zhang et al., 2023). While these approaches laid the groundwork for controllable speech generation, their dependence on fixed templates and limited attribute spaces restricts scalability toward more flexible, instruction-driven voice design. Recent studies increasingly frame controllable speech generation as an alignment problem across cross-modal representations. UniStyle (Zhu et al., 2024) unifies textual and acoustic style cues by mapping them—via a UniConnector—into a shared latent space, which is then conditioned by a VALL-E-style (Wang et al., 2023) autoregressive decoder. This design allows either a short text prompt or a 3-second reference audio to guide prosody without additional adapters. FleSpeech (Li et al., 2025) extends this idea further, explicitly aligning text, audio, and visual captions within a unified latent space. Similarly, HiStyle (Zhang et al., 2025c) adopts a hierarchical design, stacking two lightweight Transformers that sequentially convert a text prompt into coarse speaker and fine-grained style embeddings. Despite their architectural differences, these approaches all compress rich voice attributes into a single continuous vector, leaving the downstream generator with only an implicit, entangled control signal. While effective, this design imposes an inherent limitation: the control input is low-bandwidth and lacks explicit factorization of acoustic attributes, leaving fine-grained, compositional control coarse and difficult to interpret. Recent work has begun to explore direct instruction understanding via large language models to achieve more interpretable, fine-grained, and controllable speech generation. For example, VoxInstruct (Zhou et al., 2024) unifies content and style into a single natural-language instruction and employs an MT5 encoder (Xue et al., 2021) with a LLaMA-style decoder (Touvron et al., 2023) to generate discrete speech semantic tokens. Despite supporting free-form instructions, VoxInstruct compresses content and style into a single token sequence, limiting precise control over fine-grained prosodic attributes.
@@
Nevertheless, most existing systems still generate speech primarily conditioned on text and reference audio, offering limited direct control over fine-grained acoustic attributes such as pitch, speaking rate, age, emotional expression, speaking style, and others. This gap reveals a fundamental bottleneck in current generative systems: although natural language has become the dominant interface for controlling complex multimodal generation, speech synthesis still lacks a principled, flexible mechanism for translating high-level linguistic intent into fine-grained acoustic realization. In contrast to visual and video generation, voice generation remains heavily constrained by reference-based conditioning or rigid control tokens.
Recent advances in audio-centric foundation models, such as MiMo-Audio (Zhang et al., 2025a) and Step-Audio2 (Wu et al., 2025), demonstrate that scaling both model capacity and training data can substantially enhance representation quality and instruction-following capability in speech models. Their strong performance shows that large-scale models can better capture the semantics of textual instructions and generalize across speakers and styles, laying the groundwork for unified text-and-audio modeling. These developments suggest that LLM-like modeling with sufficient data and parameters is a promising approach to achieving fine-grained, natural-language-driven voice control, motivating the design of our method.
@@
To address these limitations and build on insights from recent audio-centric foundation models, we propose VoiceSculptor. This unified and highly flexible speech synthesis framework bridges natural language intent and fine-grained voice generation. Unlike conventional TTS systems that rely solely on fixed control tokens or reference audio, VoiceSculptor enables users to design speaker timbre and manipulate multiple voice attributes directly via free-form natural language instructions. By providing both open-source accessibility and instruction-driven, fine-grained controllability, our method closes a key gap between existing open-source frameworks, which often lack expressiveness, and commercial closed-source systems, which offer limited transparency and reproducibility. Our method closes a key gap between existing open-source frameworks and commercial closed-source systems.
@@ 2.1 Overview
As shown in Figure 1, VoiceSculptor adopts LLaSA-3B (Ye et al., 2025) as the voice design model and CosyVoice2 (Du et al., 2024) as the voice cloning model. LLaSA is built upon the open-source LLaMA (Touvron et al., 2023) family released by Meta and is fine-tuned to leverage the strong text understanding and sequence modeling capabilities of large language models. To enable speech generation within an LLM framework, LLaSA incorporates an advanced neural audio codec, XCodec2, that converts continuous speech waveforms into discrete audio tokens resembling text tokens.
With the introduction of XCodec2, speech synthesis is reformulated as a sequence-to-sequence generation problem. Given a natural language instruction, the LLM is responsible for interpreting the semantic and stylistic intent of the text and predicting a corresponding sequence of discrete audio tokens. These predict the audio codec, then decode tokens to reconstruct high-quality speech waveforms. This design allows the model to jointly reason over text and speech tokens using a unified autoregressive generation paradigm.
Built on this foundation, VoiceSculptor comprises three key contributions. First, we construct a comprehensive data processing pipeline that supports large-scale data collection, filtering, multi-dimensional annotation, and human verification, providing high-quality supervision for natural-language-driven voice design. Second, we introduce CoT–based fine-grained attribute modeling, which enables precise, interpretable control over prosodic and stylistic attributes by explicitly guiding the LLM to reason over attribute-related semantics during audio token generation. Third, to improve robustness and generalization to diverse, out-of-domain natural language instructions, we incorporate a RAG mechanism that provides semantically aligned in-domain instruction examples at inference time, effectively grounding the instruction interpretation process.
Together, these components enable VoiceSculptor to translate natural language instructions into controllable, reusable voice representations that can be seamlessly consumed by downstream speech synthesis models such as CosyVoice2.
@@ 2.2 Data Processing Pipeline
For linguistic alignment, automatic speech recognition (ASR) and forced alignment are applied to obtain accurate transcriptions and punctuation information. Specifically, FireRedASR (Xu et al., 2025b) is used to transcribe Chinese audio, Whisper (Radford et al., 2022) is employed for English audio, and SenseVoice (An et al., 2024) is used for ASR cross-validation, as well as language and emotion recognition. The resulting transcripts are further processed using the Kaidi  [^9] https://github.com/sunsetsonwheels/kaidi alignment tool to perform character-level (for Chinese) or word-level forced alignment, producing precise timestamps and pause durations between adjacent tokens. Based on these fine-grained temporal cues, a trained punctuation-prediction model is applied to restore punctuation, yielding linguistically coherent and temporally aligned text–audio pairs.
@@
Based on the aligned text–audio representations, a structured annotation process is subsequently conducted across multiple levels. At the advanced acoustic level, each audio sample is first analyzed using Gemini 2.5 Pro (Comanici et al., 2025) to obtain multi-dimensional annotations, including pitch, speaking rate, loudness, speaker gender and age, emotional state, paralinguistic characteristics, and contextual attributes. Based on these structured annotations, DeepSeek (Liu et al., 2025) is subsequently employed to perform translation and caption generation, producing natural-language descriptions of the audio style and vocal characteristics. To mitigate hallucinations introduced by large language models, we apply a set of rule-based regular expression filters to remove inconsistent or unsupported content. This process yields relatively high-quality, semantically grounded annotations and captions that serve as reliable supervision for natural-language-driven voice design. At the emotional level, speech is annotated through a cross-validated emotion labeling process. Specifically, we employ multiple complementary models, including Emo2Vec (Ma et al., 2024) , Qwen3-72B (Yang et al., 2025) , SenseVoice, and Qwen3-Omni (Xu et al., 2025a), to independently predict emotion-related attributes from the audio. The outputs from these models are then cross-validated to resolve inconsistencies and improve label reliability, yielding a final set of emotion annotations with greater robustness and accuracy. At the prosodic level, fine-grained acoustic attributes are annotated through a combination of automatic estimation, statistical analysis, and human verification. Specifically, we employ the DataSpeech (Lyth and King, 2024) model to estimate continuous prosodic features, including the mean and standard deviation of pitch, mean and standard deviation of energy, speaking rate, and loudness. Speaker gender and age are annotated using VoxProfile (Feng et al., 2025). Based on the extracted attributes, we perform distributional analysis across the dataset and conduct targeted human listening to calibrate attribute boundaries. The continuous prosodic features are discretized into 5 intervals per dimension, while age is categorized into 4 groups (child, youth, middle-aged, and elderly). Finally, for samples from in-house datasets, we leverage available ground-truth annotations to correct and refine the predicted age and gender labels, yielding a set of reliable, structured prosodic annotations.
@@
Through this comprehensive, iterative pipeline, VoiceSculptor constructs a high-quality, multidimensionally annotated dataset that supports natural-language-driven voice control, fine-grained attribute manipulation, and robust instruction understanding.
@@ 2.3 CoT–based Fine-grained Attribute Modeling
Various fine-grained acoustic attributes in speech signals—such as pitch, loudness, speaking rate, and temporal dynamics—play a critical role in shaping prosodic structure, expressive rhythm, and overall vocal style. However, directly conditioning speech synthesis models on explicit attribute tokens often leads to brittle control and over-reliance on structured inputs, limiting the models’ ability to generalize to diverse natural-language instructions.
@@
Moreover, to strengthen alignment between linguistic instruction and the generated speech representation, we adopt a joint cross-entropy (CE) training objective over both text and speech tokens. Specifically, the model is trained to predict not only the acoustic attribute tokens and intermediate linguistic tokens, but also the corresponding discrete speech tokens produced by the neural codec. By simultaneously optimizing CE loss across both modalities, the model learns a unified latent space that tightly couples natural-language semantics with low-level acoustic realization.
This multi-level supervision provides richer learning signals than training on speech tokens alone. In particular, the text-token CE loss encourages the model to capture semantic intent and attribute relationships better, while the speech-token CE loss guides accurate acoustic generation. Empirically, we observe that this joint optimization yields consistent performance improvements in instruction-following controllability, leading to more stable and precise attribute rendering in the synthesized speech.
Furthermore, to prevent over-reliance on explicit attribute tokens and to encourage a more profound understanding of instructions, we introduce a stochastic attribute token dropout strategy during training. Fine-grained attribute tokens are randomly removed from the input with a predefined probability, such as 0.2, forcing the model to infer the intended acoustic attributes solely from natural language instructions and contextual cues. This training strategy serves as an effective regularizer, improving the model’s robustness and generalization in natural-language-driven voice control.
@@ 2.4 Retrieval-augmented Instruction Generalization
During inference, when retrieval augmentation is enabled, the incoming natural language instruction is first converted into a dense vector representation using the same embedding model. A cosine similarity–based semantic search is then performed against the vector database to identify the most relevant in-domain instructions. The retrieved instructions, semantically aligned with the input query, are subsequently injected into the model’s input, guiding it toward a more stable and accurate interpretation of the user’s intent.
@@ 3 Experiments
We conduct extensive experiments to comprehensively evaluate the effectiveness, controllability, and scalability of the proposed VoiceSculptor framework, with a particular focus on its voice design (VD) module. Given that our training data primarily consists of Chinese instruction–speech pairs, we adopt the InstructTTSEval-Zh benchmark (Huang et al., 2025) as the primary evaluation protocol to assess instruction-following performance in a controlled and fair setting. We compare VoiceSculptor against both strong open-source baselines and representative commercial systems, and further analyze its performance under different model sizes, data scales, and training strategies. In addition, we perform a series of ablation studies to isolate the contributions of key components, including CoT-based fine-grained attribute tokens, text-side cross-entropy supervision, and RAG. Human subjective evaluations are also conducted to complement automated metrics and to validate adherence to perceptual instructions. Overall, the experimental results consistently demonstrate that VoiceSculptor achieves state-of-the-art performance among open instruction-following TTS systems, while exhibiting robust scalability and strong controllability across a wide range of settings.